OpenStudy (anonymous):

Salma has scored 84, 69, 91, 86, and 79 on her previous five tests. What score does she need on her next test so that her average (mean) is 84?

OpenStudy (anonymous):

is the answer 53 ?

OpenStudy (anonymous):

68?

OpenStudy (ikram002p):

average is sum of values then devided them on number of values , let fifth value x then average v= (84 + 69 + 91 +86+ 79+x)/5=84 solve for x

OpenStudy (anonymous):

wouldnt we divide it by 6 and not 5 because when she takes another test it will be 6 ?

OpenStudy (anonymous):

she needs an 82 ?

OpenStudy (ikram002p):

yes sorry devided by 6 :)

OpenStudy (anonymous):

so it will be 68 % ?

OpenStudy (ikram002p):

v= (84 + 69 + 91 +86+ 79+x)/6=84 409+x=6*84 409+x=504 x=504-409=95